A statistics professor drew a random sample of 81 observations and found that

x bar: 17

S 13.

Estimate the LCL of the population mean with 90% confidence.

Report your answer to two decimal places.
Answer #1

Given that,

= 17

s =13

n = 81

Degrees of freedom = df = n - 1 =81 - 1 = 80

At 90% confidence level the t is ,

= 1 - 90% = 1 - 0.90 = 0.1

= 0.1

t ,df = t0.1,80 = 1.292    ( using student t table)

Margin of error = E = t,df * (s /n)

= 1.292 * (13 / 81) = 1.8662

The 90% confidence interval estimate of the population mean is,

- E

17- 1.8662

15.13

LCL =15.13